OpenStudy (anonymous):

A gun is fired horizontally at the center of a target's bull's eye. If the target is located 100m away and the bullet is traveling 500 m/s, how far beneath the center of the bull's eye does the bullet impact?

OpenStudy (anonymous):

classic kinematic problem

OpenStudy (anonymous):

yeah, but idk where to start..

OpenStudy (anonymous):

bullet verticle equation h(t)= h_0 + v_0t + 1/2 a t^2 gravity =a=-9.8 h(t) = -4.9 t^2

OpenStudy (anonymous):

we gotta find t velocity 500m/s to travel 100m m/(m/s) = s 100/ 500 = 1/5 sec

OpenStudy (anonymous):

h(t) = -4.9 ( 1/5)^2

OpenStudy (anonymous):

.196

OpenStudy (anonymous):

so, .196 m below the bulls eye?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

Wow, that's really easy...

OpenStudy (anonymous):

Bro, your amazing.. I have a couple others ima stuck on... you mind?